OpenStudy (anonymous):

Help solve matrix. [6 2]-x=[-1 4] [-3 4] [3 -5]

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

is it x alone ??

OpenStudy (gorv):

\[\left[\begin{matrix}6 & 2 \\ -3 & 4\end{matrix}\right]-x\]

OpenStudy (gorv):

???

OpenStudy (anonymous):

Yes

OpenStudy (gorv):

ohh so x is a matrix

OpenStudy (gorv):

let it be A-X=B

OpenStudy (gorv):

now add x on both side

OpenStudy (gorv):

@-EtherealEbullience

OpenStudy (anonymous):

How do I add x?

OpenStudy (gorv):

A-X+X=B+X

OpenStudy (gorv):

what u will get..remember A is matrix we considered it as A for convinience in writing

OpenStudy (anonymous):

I'm lost. . .Can you display an example?

OpenStudy (gorv):

x-2=5 now we add 2 on both side that is left and right

OpenStudy (gorv):

x-2+2=5+2 2 will cancel -2

OpenStudy (gorv):

x=7

OpenStudy (anonymous):

x-6=-1 -6 + 6 + x = -1 + 6 x = 5?

OpenStudy (gorv):

yeah lollll..u got it

OpenStudy (gorv):

A-X+X=B+X what u will get here??

OpenStudy (anonymous):

I don't get itttttt. : I

OpenStudy (anonymous):

Too many x's I don't know what they are, confusing me.

OpenStudy (gorv):

u know to add two matric ???

OpenStudy (anonymous):

Yes, I could add them, but the x is confusing me.

OpenStudy (gorv):

let \[x=\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\]

OpenStudy (gorv):

we dont know element of matrix X so we considered it a,b,c,d ....can u subtract this from the matrix in start???

OpenStudy (anonymous):

I don't think so. o_o

OpenStudy (gorv):

u know matrix substraction??

OpenStudy (gorv):

\[\left[\begin{matrix}6 & 2 \\ -3 & 4\end{matrix}\right]-\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\]

OpenStudy (gorv):

6-a 2-b like this do the remaining 2

OpenStudy (anonymous):

-3-c 4-d. . . .Mhmmm

OpenStudy (gorv):

yeah

OpenStudy (gorv):

\[\left[\begin{matrix}6-a & 2-b \\ -3-c & 4-d\end{matrix}\right]=\left[\begin{matrix}-1 & 4 \\ 3 & -5\end{matrix}\right]\]

OpenStudy (gorv):

compare these now

OpenStudy (gorv):

corresponding element will be equal

OpenStudy (gorv):

6-a=-1 6+1-a=-1+1 7-a=0 7=a

OpenStudy (gorv):

similarly found b ,c and d

OpenStudy (anonymous):

a=7, b=-2, c=-6, d=9

OpenStudy (anonymous):

Thanks